# GAPartManip: A Large-scale Part-centric Dataset for Material-Agnostic Articulated Object Manipulation

### v1

@@ Abstract
Effectively manipulating articulated objects in household scenarios is a crucial step toward achieving general embodied artificial intelligence. Mainstream research in 3D vision has primarily focused on manipulation through depth perception and pose detection. However, in real-world environments, these methods often face challenges due to imperfect depth perception, such as with transparent lids and reflective handles. Moreover, they generally lack the diversity in part-based interactions required for flexible and adaptable manipulation. To address these challenges, we introduced a large-scale part-centric dataset for articulated object manipulation that features both photo-realistic material randomizations and detailed annotations of part-oriented, scene-level actionable interaction poses. We evaluated the effectiveness of our dataset by integrating it with several state-of-the-art methods for depth estimation and interaction pose prediction. Additionally, we proposed a novel modular framework that delivers superior and robust performance for generalizable articulated object manipulation. Our extensive experiments demonstrate that our dataset significantly improves the performance of depth perception and actionable interaction pose prediction in both simulation and real-world scenarios.
@@ I Introduction
Secondly, there is no method that can predict stable and actionable interactive poses across categories for articulated objects. Some work employs heuristic-based methods [9] to interact with articulated objects, but it is limited in diversity and fails to account for the geometric details necessary for robust interactions in real-world settings [3]. Some methods for rigid objects grasping pose prediction can generate stable poses. However, due to the lack of data on articulated objects, it is challenging to discern whether each link can interact independently, resulting in poses that are mostly non-actionable [16]. Affordance-based methods [2, 13, 17] receive widespread attention for interacting with articulated objects by generating heatmaps. However, these heatmaps are ambiguous, hard to annotate and struggle to produce stable grasping interactive poses [12].
In this paper, we address these limitations from a data-centric perspective. We introduce GAPartManip, a novel large-scale synthetic dataset that features two important aspects: (1) realistic, physics-based IR image rendering of various parts in diverse scenes, and (2) part-oriented actionable interaction pose annotations for a wide range of articulated objects. Our GAPartManip inherits 918 object instances across 19 categories from the previous GAPartNet dataset[9]. By leveraging these assets, we developed a data generation pipeline for part manipulation, producing the synthetic data needed to address the previously mentioned limitations. To improve generalizability and mitigate the sim-to-real gap, we incorporate domain randomization techniques [15] during data generation, ensuring a diverse range of outputs. In total, our dataset contains approximately 14000 scene-level samples with 8 billion part-oriented actionable pose annotations, encompassing a wide array of physical materials, object states, and camera perspectives.
Trained on the proposed dataset, we can obtain a depth reconstruction network and an actionable pose prediction network separately to address the two limitations mentioned earlier. Moreover, we compose these two neural networks modular to a novel articulated object manipulation framework. Through extensive experiments on both synthetic and real worlds, our method achieves state-of-the-art (SOTA) performance in both individual module experiments and part manipulation experiments.
@@
  • We introduce GAPartManip, a novel large-scale dataset of various articulated objects featuring realistic, physics-based rendering and diverse scene-level, part-oriented actionable interaction pose annotations.
@@ II Related work
Figure 2: Data Examples in GAPartManip.  GAPartManip is a novel large-scale synthetic dataset for articulated objects, featuring two important aspects: (1) realistic, physics-based IR rendering for various object materials in diverse scenes, and (2) part-oriented actionable interaction pose annotations for a wide range of articulated objects. Each column shows a data sample. From top to bottom, each column displays the RGB image, the IR image (only the left IR image is shown here), and the scene-level actionable interaction pose annotations.
@@ II-A Articulated object dataset
Articulated object dataset and modeling is a crucial and longstanding research field in 3D vision and robotics, encompassing a wide range of work in perception [9, 18, 19, 20, 21, 22, 23], generation [24, 25, 26, 27, 28], and manipulation [9, 29, 30, 11, 31, 10, 32, 33]. As to manipulation dataset, GAPartNet[9] annotates 6-DoF part pose to manipulate parts. Graspnet[34] and Contact-grasp[35] build several datasets, but these datasets all focus on rigid objects, neglecting the kinematic semantics specific to articulated objects. Where2act[2] first introduces a data generation pipeline for articulated objects, and they generate data by sampling successful poses in the simulator. AO-Grasp[16] leverages a curvature-based sampling method to accelerate data collection efficiency and proposes an 87k dataset of actionable poses. RPMart[12] manually annotated affordance maps for articulated objects and provided rendered data in SAPIEN[1]. None of the current datasets provide sufficient photo-realistic rendering data to improve algorithms’ perception of articulated objects during sim2real, limiting real-world performance, especially with imperfect point clouds [12, 10]. Additionally, the data collection processes are inefficient and result in small datasets, hindering algorithm generalization to unknown objects. This work aims to create a large-scale dataset with diverse photo-realistic and actionable pose data covering all types of GAParts.
@@ II-B Articulated object manipulation
Due to unique kinematic structures and geometric shapes, articulated objects present significant challenges in manipulation. Current methods can be broadly categorized into: learning-based methods and prediction-planning methods. Learning-based methods, such as reinforcement learning [10, 31] and imitation learning [32, 36], require a large amount of high-quality robot demonstration. However, collecting such data is both impractical and time-consuming, and their sim-to-real performance heavily relies on simulator. Current prediction-planning methods[2, 9, 11, 37, 38, 39] focus on visual affordance but offer ambiguous interactive poses and struggle to generalize due to limited data. They rely on 3D point clouds, ignoring the impact of object materials. In the real world, depth cameras often miss critical points like handles and lids, reducing sim-to-real performance.
@@ III-A Overview
We build a photo-realistic rendering pipeline for each asset in indoor scenes. We render RGB images, IR images, depth maps, and part-level segmentations. Additionally, we create high-quality and physics-plausible interaction pose annotations for each part of the articulated object. Then leverage our GPU-accelerated scene-level pose annotation pipeline to generate dense, part-oriented actionable interaction pose annotations for each data sample. Our dataset contains over 8 billion actionable poses across 241680 data samples. Fig. 2 shows examples of data samples from our dataset. Our whole data generation pipeline is illustrated in Fig. 3.
Figure 3: Dataset Generation Pipeline. For scene-level data sample rendering, we input the object asset into our photo-realistic rendering pipeline, generating one RGB image and two IR images (left and right) for each camera perspective. For pose annotation, we begin by performing mesh fusion on each GAPart of the object to establish a one-to-one correspondence between GAParts and meshes. Then, we use FPS to obtain the point cloud for each GAPart, enabling part-level stable interaction pose annotation. These poses are further utilized for scene-level actionable interaction pose annotation for each rendered data sample.
@@ III-B Photo-realistic Scene-level Rendering
Our photo-realistic rendering pipeline is built upon NVIDIA Isaac Sim [40]. Specifically, we simulate the RGB and IR imaging process of Intel RealSense D415, a widely-used structured light camera for real-world depth estimation in previous research works. We replicate the layout of the D415 imaging system consisting of four hardware modules, i.e., an IR projector, an RGB camera, and two infrared (IR) cameras. We also project a similar shadow pattern onto the scenes with D415.
Inspired by previous works [41, 14], we incorporate domain randomization techniques into our rendering pipeline to mimic the IR rendering under various lighting conditions and material properties in the real world. We render each object in 20 different scenes with various domain randomization settings. Concretely, we randomly vary ambient lighting, background, and object material properties in the scene, generating more diverse data that covers a wider range of real-world imaging conditions. we further randomize the ambient light positions and intensities within each scene. More importantly, we randomize the parameters of all diffuse, transparent, specular, and metal materials of each part corresponding to their semantics. Finally, we uniformly randomize the joint poses of the object within its joint limits in each scene during the rendering process.
We render the objects and parts from different distances. We render each scene with 5 object-centric camera perspectives for the whole object and 5 part-centric camera perspectives for each part. To place the object within the camera view, i.e., the object-centric perspective, the camera is positioned at a latitude of ranged in [10°,60°] and a longitude ranged in [-60°,60°] in the target object. To capture the more fine-grained parts, i.e., the part-centric perspective, we leverage part pose annotations from GAPartNet and the current joint poses to determine the position and orientation of each part in the scene. The camera is then randomly positioned around each part, aiming directly toward the part center. As a result, the target part occupies the primary area of the image. During this process, camera viewpoints are randomly sampled within a latitude range of [0°,60°] and a longitude range of [-75°,75°].
@@ III-C GPU-accelerated Scene-level Pose Annotation
We employ a pose sampling strategy similar to GraspNet [34] to annotate dense and diverse stable interaction poses for each GAPart, based on the original semantic annotations in GAPartNet [9]. First, we perform mesh fusion for each part, merging the meshes corresponding to the same part to establish a one-to-one correspondence between parts and meshes. Then, we apply Farthest Point Sampling (FPS) to downsample the mesh for each part, resulting in $N$ candidate points for pose sampling. For each candidate point, we uniformly generate $V\times A\times D$ candidate poses, where $V$ is the number of gripper views distributed uniformly over a spherical surface, $A$ represents the number of in-plane gripper rotations, and $D$ refers to the number of gripper depths. In our case, $N=512$, $V=64$, $A=12$, and $D=4$. We follow GraspNet to calculate the pose score based on antipodal analysis.
@@
To obtain part-centric grasping poses, We first project the part-level grasping poses into the scene using the part pose annotations, and then filter out unreasonable and unreachable poses. More concretely, We classify grasping poses that do not align with single-view partial point clouds as unreasonable. Meanwhile, we consider poses that cause collisions between the gripper and other objects as unreachable.
However, such a filtering process is computationally demanding due to the large amounts of points in the scene. To accelerate the pose annotation process, we implemented a CUDA-based optimization for the filtering process. Our optimization significantly reduces the processing time from 5 minutes to less than 2 seconds for each part which is nearly a 150 times speed-up. As a result, the originally year-long pose reduction process can now be completed within 3 days.
@@ IV Framework
Figure 4: Framework overview. Given IR images and raw depth, the depth reconstruction module first performs depth recovery. Subsequently, the pose prediction module generates a 7-DOF actionable pose and a 3-DOF motion directive based on the reconstructed depth. Finally, the local planner module carries out the action execution.
@@ IV-A Depth reconstruction module
The input to our system is a single view RGB-D observation including a raw depth $I_{d}$, left IR image $I^{l}_{ir}$, right IR image $I^{r}_{ir}$, and an RGB image $I_{c}$. However, the raw sensor depth are often imcomplete and even incorrect because transparent and reflective surfaces are inherently ambiguous for structured light and Time-of-Flight depth cameras. Therefore, we leverage diffusion model-based approaches to estimate and restore the incomplete depths of raw sensor outputs. We use D3RoMa [14] as our depth predictor and fine-tune it on our dataset.
@@ IV-B Pose prediction module
Different from 6 DoF grapsing pose prediction for rigid object manipulation, we need to predict both the 6-DoF part grasping pose and the 2-DoF movement direction after grasping. We adapt the SOTA method Economicgrasp [42] as our actionable pose estimator dubbed Part-aware EcoGrasp and use pretrained GAPartNet[9] to predict the part movement direction.
@@
$\scriptsize s^{P}_{i}=\frac{1}{\textstyle\displaystyle\sum_{j,k}\left|A^{i,j}_{k}\right|}{c_{a}^{i}\textstyle\displaystyle\sum_{j,k}\textbf{1}\left(q_{k}^{i,j}>T\right)c_{k}^{i,j}},$ (1)
$\scriptsize s^{V}_{i}=\frac{1}{\textstyle\displaystyle\sum_{k}\left|A_{k}^{i,j}\right|}{c_{a}^{i}\textstyle\displaystyle\sum_{k}\textbf{1}\left(q_{k}^{i,j}>T\right)c_{k}^{i,j}},$ (2)
where $T$ is a predefined threshold to filter out inferior quality poses. We then train Part-aware EcoGrasp[42] following [42]. Additionally, we utilize the pre-trained GAPartNet [9] to predict the motion direction which speficies the part movement direction after grasping the actionable part.
@@ IV-C Local planner module
We use CuRobo [43] as our motion planner. It optimizes motion trajectories based on actionable poses given by the pose prediction module, computes joint angles through inverse kinematics, and drives the robot to execute trajectory actions through joint control modes. Subsequently, the robot executes actions based on the motion direction $\vec{r_{p}}$
@@ V-A Depth Estimation Experiments
In this section, we evaluate different depth estimation methods with our GAPartManip  to demonstrate the effectiveness of our dataset for improving articulated object depth estimation in both simulation and the real world.
Data Preparation. We split the dataset into training and test sets using an approximate 8:2 ratio. To maintain comprehensive coverage, each object category is split carefully, ensuring that both the training and test sets include samples from all categories. Additionally, we make sure that samples rendered from the same object category are assigned exclusively to either the training or test set. We compare our method with following baselines:
@@
* indicates that the method is trained on the GAPartManip dataset.
Figure 5: Qualitative Results for Depth Estimation in the Real World. Our refined depth is more robust for transparent and translucent lids and small handles compared to RAFT-Stereo. Zoom in to better observe small parts like handles and knobs.
Results and Analysis. The quantitative results in simulation are presented in Tab. I. The results indicate that the traditional stereo matching algorithm, SGM, struggles in scenes with articulated objects with challenging material characteristics. The same observation applies to the pre-trained RAFT-Stereo. Meanwhile, the pre-trained D3RoMa models demonstrate reasonably good stereo depth estimation capabilities in the experiments. However, both RAFT-Stereo and D3RoMa are significantly enhanced when fine-tuned on GAPartManip. Specifically, RAFT-Stereo achieves a 150% improvement in MAE compared to its pre-trained version, while our model exhibits a 600% improvement, achieving the best performance in the simulation. As illustrated in Fig. 5, the fine-tuned models also demonstrate strong depth estimation performance in real-world scenarios. In particular, in real-world environments with challenging materials, as shown in the first three rows of the figure, our model significantly outperforms the fine-tuned RAFT-Stereo and the raw depth, exhibiting noticeably better robustness. Both simulation and real-world experiments demonstrate the effectiveness of our proposed GAPartManip  in substantially improving depth estimation for articulated objects with challenging materials.

### v2

@@ Abstract
Effectively manipulating articulated objects in household scenarios is a crucial step toward achieving general embodied artificial intelligence. Mainstream research in 3D vision has primarily focused on manipulation through depth perception and pose detection. However, in real-world environments, these methods often face challenges due to imperfect depth perception, such as with transparent lids and reflective handles. Moreover, they generally lack the diversity in part-based interactions required for flexible and adaptable manipulation. To address these challenges, we introduced a large-scale part-centric dataset for articulated object manipulation that features both photo-realistic material randomization and detailed annotations of part-oriented, scene-level actionable interaction poses. We evaluated the effectiveness of our dataset by integrating it with several state-of-the-art methods for depth estimation and interaction pose prediction. Additionally, we proposed a novel modular framework that delivers superior and robust performance for generalizable articulated object manipulation. Our extensive experiments demonstrate that our dataset significantly improves the performance of depth perception and actionable interaction pose prediction in both simulation and real-world scenarios. More information and demos can be found at: https://pku-epic.github.io/GAPartManip/.
@@ I Introduction
Secondly, there is no method that can predict stable and actionable interaction poses across categories for articulated objects. Some work employs heuristic-based methods [9] to interact with articulated objects, but it is limited in diversity and fails to account for the geometric details necessary for robust interactions in real-world settings [3]. Some methods for rigid object grasping pose prediction can generate stable poses. However, due to the lack of data on articulated objects, it is challenging to discern whether each link can be interacted with independently, resulting in poses that are mostly non-actionable [16]. Affordance-based methods [2, 13, 17] receive widespread attention for interacting with articulated objects by generating heatmaps. However, these heatmaps are ambiguous, hard to annotate, and struggle to produce stable interaction poses [12].
In this paper, we address these limitations from a data-centric perspective. We introduce GAPartManip, a novel large-scale synthetic dataset that features two important aspects: 1) realistic, physics-based IR image rendering for various parts in diverse scenes, and 2) part-oriented actionable interaction pose annotations for a wide range of articulated objects. Our GAPartManip inherits 918 object instances across 19 categories from the previous GAPartNet dataset[9]. By leveraging these assets, we develop a novel data generation pipeline for part manipulation, producing the synthetic data needed to address the previously mentioned limitations. To improve generalizability and mitigate the sim-to-real gap, we incorporate domain randomization techniques [15] during data generation, ensuring a diverse range of outputs. In total, our dataset contains approximately 14,000 scene-level samples with 8 billion part-oriented actionable pose annotations, encompassing a wide array of physical materials, object states, and camera perspectives.
Through training on the proposed dataset, we obtain a depth reconstruction network and an actionable pose prediction network separately to address the two limitations mentioned earlier. Moreover, we compose these two neural networks modular to a novel articulated object manipulation framework. Through extensive experiments in both simulation and the real world, our method achieves state-of-the-art (SOTA) performance in both individual modular experiments and articulated object manipulation experiments.
@@
  • We introduce GAPartManip, a novel large-scale dataset with various articulated objects featuring realistic, physics-based rendering and diverse scene-level, part-oriented actionable interaction pose annotations.
@@ II Related work
Figure 2: Data Examples in GAPartManip.  GAPartManip is a novel large-scale synthetic dataset for articulated objects, featuring two important aspects: 1) realistic, physics-based IR rendering for various object materials in diverse scenes, and 2) part-oriented actionable interaction pose annotations for a wide range of articulated objects. Each column shows a data sample. From top to bottom, each row displays the RGB image, the IR image (only the left IR image is shown here), and the scene-level actionable interaction pose annotations.
@@ II-A Articulated Object Dataset
Articulated object dataset and modeling is a crucial and longstanding research field in 3D vision and robotics, encompassing a wide range of work in perception [9, 18, 19, 20, 21, 22, 23], generation [24, 25, 26, 27, 28], and manipulation [9, 29, 30, 11, 31, 10, 32, kuang2024ram]. As to manipulation dataset, GAPartNet[9] annotates 7-DoF part pose to manipulate parts. GraspNet[33] and Contact-Grasp[34] build several datasets, but these datasets all focus on rigid objects, neglecting the kinematic semantics specific to articulated objects. Where2Act[2] first introduces a data generation pipeline for articulated objects, which generates data by sampling successful poses in the simulator. AO-Grasp[16] leverages a curvature-based sampling method to accelerate data collection efficiency and proposes an 87K dataset for actionable poses. RPMart[12] manually annotates affordance maps for articulated objects and provides rendering data in SAPIEN[1]. None of the current datasets provide sufficient photo-realistic rendering data to improve the algorithm’s capability of perception for articulated objects during sim-to-real, limiting the real-world performance, especially with imperfect point clouds [12, 10]. Additionally, the data collection processes are inefficient and result in small datasets, hindering the algorithm’s generalizability to unseen objects. This work aims to create a large-scale part-centric dataset with diverse photo-realistic rendering and extensive actionable pose data for articulated object manipulation.
@@ II-B Articulated Object Manipulation
Due to complex kinematic structures and geometric shapes, articulated objects present significant challenges in manipulation. Current methods can be broadly categorized into learning-based methods and prediction-planning methods. Learning-based methods, such as reinforcement learning [10, 31] and imitation learning [32, 35], require either the realistic and accurate simulation or a large amount of high-quality robot demonstration. However, collecting such data is both impractical and time-consuming, and their sim-to-real performance heavily relies on the quality of the simulation. Current prediction-planning methods[2, 9, 11, 36, 37, kuang2024ram] focus on visual affordance but offer ambiguous interaction poses and struggle to generalize due to the limited data. These methods typically rely on accurate 3D point cloud input, ignoring the impact of object materials on the quality of point clouds. In practice, depth sensors often struggle with challenging materials like glass or metal, missing critical geometric structures on the objects such as handles and lids, significantly limiting the model’s sim-to-real performance.
@@ III-A Overview
We build a photo-realistic rendering pipeline for each asset in indoor scenes. We render RGB images, IR images, depth maps, and part-level segmentations. Additionally, we create high-quality and physics-plausible interaction pose annotations for each part on the articulated object. Then, we leverage our GPU-accelerated scene-level pose annotation pipeline to generate dense, part-oriented, actionable interaction pose annotations for each rendering data sample. Our dataset contains over 8 billion actionable poses across 241,680 rendering data samples. Fig. 2 shows examples of data samples from our dataset. Our whole data generation pipeline is illustrated in Fig. 3.
Figure 3: Dataset Generation Pipeline. For scene-level data sample rendering, we input the object asset into our photo-realistic rendering pipeline, generating one RGB image and two IR images (left and right) for each camera perspective. For pose annotation, we begin by performing mesh fusion on each GAPart on the object to establish a one-to-one correspondence between GAParts and meshes. Then, we use FPS to obtain the point cloud for each GAPart, enabling part-level stable interaction pose annotation. These poses are further utilized in scene-level actionable interaction pose annotation for each rendering data sample.
@@ III-B Photo-realistic Scene-level Rendering
Our photo-realistic rendering pipeline is built upon NVIDIA Isaac Sim [38]. Specifically, we simulate the RGB and IR imaging process of Intel RealSense D415, a widely-used structured light camera for real-world depth estimation in previous research works. We replicate the layout of the D415 imaging system consisting of four hardware modules, i.e., an infrared (IR) projector, an RGB camera, and two IR cameras. We also project a similar shadow pattern onto the scenes with D415.
Inspired by previous works [39, 14], we incorporate domain randomization techniques into our rendering pipeline to mimic the IR rendering under various lighting conditions and material properties in the real world. We render each object in 20 different scenes with various domain randomization settings. Concretely, we randomly vary ambient lighting, background, and object material properties in the scene, generating more diverse data that covers a wider range of real-world imaging conditions. we further randomize the ambient light positions and intensities within each scene. More importantly, we randomize the parameters of all diffuse, transparent, specular, and metal materials for each part corresponding to their semantics. Finally, we uniformly randomize the joint poses of the object within its joint limits in each scene during the rendering process.
We render the objects and parts from different distances. We render each scene with 5 object-centric camera perspectives for the whole object and 5 part-centric camera perspectives for each part. To place the object within the camera view, i.e., the object-centric perspective, the camera is positioned at a latitude of ranged in [10°,60°] and a longitude ranged in [-60°,60°] in the target object frame. To capture more fine-grained details of the parts, i.e., the part-centric perspective, we leverage part pose annotations in GAPartNet and the current joint poses to determine the position and orientation of each part in the scene. The camera is then randomly positioned around each part, aiming directly toward the part center. As a result, the target part occupies the primary area of the image. During this process, camera viewpoints are randomly sampled within a latitude range of [0°,60°] and a longitude range of [-75°,75°].
@@ III-C GPU-accelerated Scene-level Pose Annotation
We employ a pose sampling strategy similar to GraspNet [33] to annotate dense and diverse stable interaction poses for each GAPart, based on the original semantic annotations in GAPartNet [9]. First, we perform mesh fusion for each part, merging the meshes corresponding to the same part to establish a one-to-one correspondence between parts and meshes. Then, we apply Farthest Point Sampling (FPS) to downsample the mesh of each part, resulting in $N$ candidate points for pose sampling. For each candidate point, we uniformly generate $V\times A\times D$ candidate poses, where $V$ is the number of gripper views distributed uniformly over a spherical surface, $A$ represents the number of in-plane gripper rotations, and $D$ refers to the number of gripper depths. In our case, $N=512$, $V=64$, $A=12$, and $D=4$. We follow GraspNet to calculate the pose score based on the antipodal analysis.
@@
To obtain part-centric interaction poses, We first project the part-level interaction poses into the scene using the part pose annotations, and then filter out unreasonable and unreachable poses. More concretely, we classify poses that do not align with single-view partial point clouds as unreasonable. Meanwhile, we consider poses that cause collisions between the gripper and other parts of the object or the scene as unreachable.
However, such a filtering process is extremely computationally demanding due to the large amounts of points in the scene. To accelerate the pose annotation, we implement a CUDA-based optimization for the filtering process. Our optimization significantly reduces the processing time from 5 minutes to less than 2 seconds for each part, which is nearly a 150-times speed-up. As a result, the originally year-long pose annotation process can now be completed within 3 days.
@@ IV Framework
Figure 4: Framework Overview. Given IR images and raw depth map, the depth reconstruction module first performs depth recovery. Subsequently, the pose prediction module generates a 7-DoF actionable pose and a 6-DoF post-grasping motion for interaction based on the reconstructed depth. Finally, the local planner module carries out the action planning and execution.
@@ IV-A Depth Reconstruction Module
The input to our system is a single view RGB-D observation including a raw depth $I_{d}$, a left IR image $I^{l}_{ir}$, a right IR image $I^{r}_{ir}$, and an RGB image $I_{c}$. The raw sensor depth is often incomplete and even incorrect because transparent and reflective surfaces are inherently ambiguous for structured lights and time-of-flight depth sensors. To tackle this problem, we leverage diffusion model-based approaches to estimate and restore the incomplete depth for raw sensor outputs. Specifically, we use D3RoMa [14] as our depth predictor and fine-tune it on our dataset.
@@ IV-B Pose Prediction Module
Different from the 7-DoF grasping pose prediction for rigid object manipulation, here we need to predict both the 7-DoF actionable interaction pose and the 6-DoF interaction motion.
@@
$s^{P}_{i}=\frac{1}{\textstyle\displaystyle\sum_{j,k}\left|A^{i,j}_{k}\right|}{c_{act}^{i}\textstyle\displaystyle\sum_{j,k}\textbf{1}\left(q_{k}^{i,j}>T\right)c_{k}^{i,j}},$ (1)
$s^{V}_{i,j}=\frac{1}{\textstyle\displaystyle\sum_{k}\left|A_{k}^{i,j}\right|}{c_{act}^{i}\textstyle\displaystyle\sum_{k}\textbf{1}\left(q_{k}^{i,j}>T\right)c_{k}^{i,j}},$ (2)
where $T$ is a pre-defined threshold to filter out inferior-quality poses. We train the Part-aware EcoGrasp following the original setting [40].
We utilize the pre-trained GAPartNet [9] to predict the interaction motion, which specifies the post-grasping movement of the end-effector after grasping the actionable part for interaction.
@@ IV-C Local Planner Module
We use CuRobo [41] as our motion planner. The planner optimizes motion trajectories to the actionable poses estimated by the pose prediction module, computes robot joint angles through inverse kinematics, and drives the robot to execute trajectory actions through joint control. After reaching the target pose, the robot executes grasping for interaction and subsequently executes the post-grasping motion from the pose prediction module to complete the interaction.
@@ V-A Depth Estimation Experiments
In this section, we evaluate different depth estimation methods with our GAPartManip dataset to demonstrate the effectiveness of our dataset for improving articulated object depth estimation in both simulation and the real world.
Data Preparation. We split the dataset into training and testing sets using an approximate 8:2 ratio. To maintain comprehensive coverage, each object category is split carefully, ensuring that both the training and testing sets include samples from all categories. Additionally, we make sure that samples rendered from the same object category are assigned exclusively to either the training or testing set. We compare our method with following baselines:
@@
* indicates that the method is fine-tuned on the GAPartManip dataset.
Figure 5: Qualitative Results for Depth Estimation in the Real World. Our refined depth maps are cleaner and more accurate than the ones from the baseline, indicating that our depth reconstruction module is more robust for transparent and translucent lids and small handles. Zoom in to better observe small parts like handles and knobs.
Results and Analysis. The quantitative results in simulation are presented in Tab. I. The results indicate that the traditional stereo matching algorithm, SGM, struggles in scenes with articulated objects with challenging material characteristics. The same observation applies to the pre-trained RAFT-Stereo. Meanwhile, the pre-trained D3RoMa models demonstrate reasonably good stereo depth estimation capabilities in the experiments. However, both RAFT-Stereo and D3RoMa are significantly enhanced when fine-tuned on GAPartManip. Specifically, RAFT-Stereo achieves a 150% improvement in MAE compared to its pre-trained version, while our model exhibits a 600% improvement in MAE, achieving the best performance in the simulation. As illustrated in Fig. 5, the fine-tuned models also demonstrate strong depth estimation performance in real-world scenarios. In particular, in real-world environments with challenging materials, as shown in the first three rows of the figure, our model significantly outperforms the fine-tuned RAFT-Stereo and the raw depth, exhibiting noticeably better robustness. Both simulation and real-world experiments demonstrate the effectiveness of our proposed GAPartManipin substantially improving depth estimation for articulated objects with challenging materials.